# Going beyond density functional theory accuracy: Leveraging experimental data to refine pre-trained machine learning interatomic potentials

### v1

@@ Results
To achieve beyond DFT-level accuracy, we refine the pre-trained MLIP of uranium dioxide and uranium mononitride to match target synthetic EXAFS spectra. The unrefined MLIPs are pre-trained to DFT/DFT+U total energy and atomic forces, while the refined MLIPs are re-trained with transfer learning to the shifted atomic energies. Although only atomic-scale quantities are involved in the initial training and refinement process, the refined MLIP predicts improved macroscopic properties, beyond EXAFS, that are not explicitly targeted in the training process. To validate the improvement in the refined UO2 and UN MLIPs, we present a set of benchmarks and compare the MLIP predictions with experimental data for thermophysical properties (lattice parameters, bulk modulus and heat capacity), crystal properties (elastic constants and phonon dispersion), energetic properties (point defect energies), and transport properties (diffusion coefficients).
We first verify the accuracy of the UO2 and UN unrefined MLIPs pre-trained to the respective DFT+U and DFT datasets. TableΒ 1 reports the energy-per-atom and force root-mean-square error (RMSE) between the unrefined MLIP-predicted values and the corresponding DFT+U/DFT reference values for UO2 and UN. The energy-per-atom RMSE values are similar for both MLIPs and less than 0.2 kcal/mol. The force RMSE for the UO2 MLIP is approximately 10 kcal/mol/Γ , while the force RMSE for the UN MLIP is around 3 kcal/mol/Γ . The relatively larger force RMSE value of UO2 is a consequence of the reduced DFT+U dataset consisting of more challenging atomic configurations, with fewer near-equilibrium low-energy configurations. For comparison, the RMSEs for UN (0.09 kcal/mol for energy-per-atom and 3.14 kcal/mol/Γ for forces) are only slightly higher than those reported previously (0.06 kcal/mol for energy-per-atom and 2.6 kcal/mol/Γ for forces) for the same HIP-NN architecture trained on the same dataset.Β [29] Furthermore, the RMSEs for both UO2 and UN are of the same magnitude to the reported values for a different MLIP architecture (0.1 kcal/mol for energy-per-atom and 8.8 kcal/mol/Γ for forces for UO2 and 0.03 kcal/mol for energy-per-atom and 1.7 kcal/mol/Γ for forces for UN), demonstrating that our model achieves the expected level of accuracy on these training datasets.[27, 29]
| System | energy-per-atom (kcal/mol) | force (kcal/mol/Γ ) |
| UO2 | 0.15 | 10.04 |
| UN | 0.09 | 3.14 |
@@
Since the refinement process is performed with respect to the synthetic target EXAFS spectra for both U and O, we verify that the refined MLIP predicts spectra in closer agreement with the experimental EXAFS spectrum for U. Because refinement is performed at a single temperature of 75Β K, we also verify the improvement for EXAFS spectrum prediction at 300Β K.
@@
FigureΒ 2 presents a comparison of the experimental EXAFS spectrum for uranium in UO2 with those computed using unrefined and refined MLIPs at 77Β K (a) and at 300Β K (b). At both temperatures, the spectral peaks shift towards higher k-values, aligning more closely with the experimental data. At 77Β K, the peak amplitude of the refined MLIP increases slightly compared to the unrefined MLIP. However, both the unrefined and refined spectra have an enhanced peak amplitude compared to the experimental spectra. At 300Β K, we observe better agreement of the peak amplitude with the experimental spectrum.
@@
Even though the amplitudes for the synthetic perfect crystal spectrum are significantly larger than the experimental spectrum (Figure S5), the refined MLIP predicts spectral peak amplitudes that are in reasonable agreement with experiment. Overfitting to the synthetic spectra could induce unusually strong interatomic forces, reducing thermal vibrations and producing overly large peak amplitudes. Careful transfer learning (see Simulation details section) helps avoid overfitting and catastrophic forgetting of the initial DFT+U training dataset. Figure S8 also shows systematic and gradual improvement in the MLIP-generated EXAFS spectra with respect to the synthetic target spectra, suggesting that the EXAFS refinement approach converges smoothly.
@@
TableΒ 2 reports the UO2 elastic constants and bulk modulus at 0Β K for the unrefined and refined MLIPs compared to the experimental values at 296 K[45]. The refined MLIP shows a significant improvement in the underestimated C11 and C44 constants, while only a small improvement in the overestimated C12 constant. The bulk modulus is overestimated by approximately 6% and 11% for the unrefined and refined MLIPs, respectively. This overestimation for both MLIPs is likely attributed to the temperature difference between our 0Β K calculation and experimental measurements at 296Β K.
| Experiment | 389.3 1.7 | 118.7 1.7 | 59.7 0.3 | 209.0 |
| Unrefined MLIP | 353.89 | 156.18 | 37.41 | 221.42 |
| Refined MLIP | 381.59 | 154.59 | 47.90 | 232.30 |
@@
FigureΒ 3 shows the temperature dependence for (a) lattice parameter, (b) heat capacity, and (c) bulk modulus computed using the unrefined and refined MLIPs in comparison to experimental data. The systematic and gradual improvement of these properties at each refinement iteration is shown in Figure S9. As shown in FigureΒ 3(a), the most significant improvement achieved through MLIP refinement is the enhanced accuracy of lattice parameter with respect to temperature. Because EXAFS spectra depend strongly on the local atomic environment and the interatomic distances around the absorbing atom, the improvement in lattice parameter can be attributed to a better agreement with the target EXAFS spectra, which was generated with the experimental lattice parameter. Remarkably, the refined MLIP produces improved lattice parameters over a range of temperatures from 300 K to 3000 K, even though the refinement is performed at a low temperature of 75 K. FigureΒ 3(a) demonstrates that the nearly constant overprediction of lattice parameter by the unrefined MLIP gradually deflects at higher temperatures above 2000 K. This deflection is also improved in the lattice parameters predicted by the refined MLIP, suggesting that the refined MLIP is more stable at higher temperatures. This improvement can be more clearly observed from the thermal expansion plot at higher temperature from 2500 K to 3000 K (Figure S10). Such consistent improvement over a range of temperature is not observed when the re-training is done without freezing certain HIP-NN layers (Figure S11 and Figure S12), which implies that freezing some layers during transfer learning eliminates possible overfitting to the small refinement dataset.
@@
FigureΒ 3(b) shows the temperature-dependent heat capacity for unrefined and refined MLIPs compared to the experimental results. The increase in heat capacity with temperature for UO2 has been attributed to different factors in different temperature ranges. At low temperatures up to 1000 K, the heat capacity increases because of the harmonic lattice vibrations. In the temperature range from 1000 K to 1500 K, there is an increase in anharmonicity. From 1500 K to 2670 K, the increase in heat capacity is due to the formation of lattice and electronic defects[46]. The abrupt transition in heat capacity observed experimentally at 2670 K, i.e., the Bredig temperature, is attributed to the saturation of defect concentration[49]. While both refined and unrefined MLIPs reproduce this transition phenomenon, the refined MLIP predicts a slightly more accurate transition temperature compared to the unrefined MLIP, again indicative of a better description of UO2 at high temperatures.
@@
FigureΒ 3(c) shows improvement in the refined MLIP for bulk modulus temperature dependence. The predicted bulk modulus using unrefined MLIP is underestimated with respect to the experimental results, except at a relatively low temperature of 300 K. In contrast, the refined MLIP predicts bulk modulus values that are consistent over the entire temperature range and closer in agreement with the experimental values, especially at higher temperatures. At 300 K, the refined MLIP overestimates the bulk modulus compared to the values from Hutchings et al. [48] However, the refined MLIP predicts a bulk modulus value at room temperature in close agreement with the experimental data reported by Fritz et al. [45] at 296 K as shown in TableΒ 2. Overall, the refined MLIP predicts higher bulk modulus compared to the unrefined MLIP, consistent with higher EXAFS peak amplitudes and a stiffer response to volumetric compression.
@@
FigureΒ 4 shows that the phonon dispersion spectra are similar for both the unrefined and refined MLIP. This similarity shows that the UO2 vibrations are not affected significantly after refinement. The synthetic target spectra is computed for a perfect crystal. Overfitting to the synthetic target spectra during refinement would lead to stiffer vibrations and worse phonon spectra. As the phonon spectra remain relatively unchanged, the refinement improves the structural properties while not causing the refined MLIP to forget the vibrational information contained in the DFT+U original training dataset.
Nuclear fuels constantly undergo neutron irradiation, which results in vacancies, interstitial and dislocation defects. The defect energies are an important property for a nuclear fuel as the defect behavior helps determine mechanical degradation and fuel operation lifetime. The unrefined MLIP predicts reliable defects in comparison with DFT+U. Because the refinement process shifts the atomic energies for the MLIP, it is important to validate that the defect energies remain reasonable. TableΒ 3 presents the point defect energies calculated for the unrefined and refined MLIPs compared to the experimental and DFT+U literature values.[51, 52, 53] The Schottky defect energy calculated using the unrefined and refined MLIPs lies within the experimental range. The Frenkel pair defect energies for U and O for both the unrefined and refined MLIPs fall in the DFT+U energy range, but are slightly overestimated compared to the experimental values. Considering the wide uncertainty ranges in the experimental and DFT+U values, both the unrefined and refined MLIPs predict reasonable defect energies, demonstrating that the refinement process does not negatively impact defect energies.
@@
Validation of transport properties is especially important when training to only static structural data, such as EXAFS spectra. For this reason, we compare the diffusion coefficient for oxygen in UO2 predicted with the unrefined and refined MLIPs. The technical details for calculating the diffusion coefficients are provided in the Simulation Details section and Supporting Information (Figure S13). Figure 5 shows the logarithm of the oxygen diffusion coefficient as a function of inverse temperature (1000/T) (K-1) from 2100 K to 3000 K. The diffusion coefficient increases by around an order of magnitude after refinement. As experimental diffusion coefficient data are not available for validation, we compare with experimental activation energies for oxygen diffusion in UO2. FigureΒ 5 shows clearly that there are two regions of linear dependency. The transition temperatures for the unrefined MLIP (2600 K) and refined MLIP (2700 K) coincide closely with the experimental Bredig temperature of 2670 K. The activation energy is calculated separately for the low-temperature and high-temperature regions.
TableΒ 4 reports the activation energy of oxygen diffusion in UO2 calculated for the unrefined and refined MLIPs. At high temperatures, the activation energy agrees closely with the experimental values. The activation energy at low temperatures is overestimated for both unrefined and refined MLIPs. Notably, the re-training has minimal impact on the activation energy in both ranges of temperature, demonstrating that the MLIP has retained valuable transport information from the DFT+U dataset.
@@
To further validate our methodology, we refine an MLIP for uranium mononitride that was pre-trained to DFT. data[29] In contrast to UO2, DFT+U calculations were not required to properly describe the magnetic and metallic character of UN. Therefore, the unrefined UN MLIP predictions do not differ as strongly from experiment, especially for the lattice parameter. Consequently, unlike UO2, the EXAFS peak positions obtained from the unrefined MLIP already align well with the synthetic target EXAFS spectra for UN (Figure S14). Although the improvement needed for the UN system is less than for UO2, it is still important to carefully refine the MLIP without depreciating its accuracy to predict other properties that it has learned well from the DFT training dataset.
@@
Because the refinement procedure uses synthetic target spectra for U and N at a single low temperature of 50Β K, we verify that the agreement also improves with the experimental U EXAFS spectra, even at higher temperatures. FigureΒ 6 shows the comparison of the U EXAFS spectra for unrefined and refined MLIPs with respect to the experimental EXAFS spectra at 80Β K (a) and 300Β K (b). At a high level, the unrefined and refined MLIPs reproduce almost every broad oscillation present in the experimental EXAFS spectra. In a closer look, the EXAFS peak positions of the refined MLIP align slightly better with the experimental results compared to the unrefined MLIP. However, the simulated spectra are smooth and do not capture certain sharp features/peaks in the experimental spectra. We attribute this smoothness to the cutoff radius of 6 Γ used to calculate the simulated spectra. Increasing the cutoff can capture more nuanced structural information, resulting in more detailed EXAFS spectra[55, 56]. However, a greater cutoff radius increases the number of scattering paths from distant coordination shells, which dramatically increases the computational cost of calculating the EXAFS spectra. Selecting a 6 Γ cutoff radius is adequate for resolving key structural features in EXAFS while keeping computational costs manageable during the refinement process.
@@
Similar to UO2, the synthetic target EXAFS spectrum of U in UN has a larger amplitude compared to the experimental spectra. The synthetic spectra are calculated for a perfect crystal structure, which eliminates the dampening of peaks due to atomic vibrations. Despite the larger amplitude of target spectra, the refined MLIP actually predicts a slight reduction in peak amplitude (Figure S14), which improves agreement with the experimental EXAFS spectrum (FigureΒ 6). The refinement also causes a slight leftward shift in both the U and N EXAFS spectra, mainly in the range of the wavenumber k = 8 to 15. The leftward shift of the EXAFS peaks leads to an increase in the interatomic distances, signifying a decrease in the interatomic forces. With more flexible vibrations, the mean-square-displacement of UN atoms increases, which slightly dampens the spectra. The evidence of weaker interatomic forces is supported by certain properties used to validate the refined MLIP, discussed below.
@@
To validate the refined MLIP for UN against structural properties, TableΒ 5 reports the elastic tensors and the bulk modulus at 0Β K using the unrefined and refined MLIPs as well as the experimental values at 290 K.[57] The values for C12 and C44 improve, while C11 is further underestimated after refinement. The bulk modulus prediction improves significantly. The large drop in the bulk modulus on refinement is consistent with a reduction in interatomic forces.
| Experiment | 423.9 0.6 | 98.1 0.9 | 75.7 0.2 | 206.7 0.8 |
| Unrefined MLIP | 387.97 | 131.94 | 49.78 | 219.48 |
| Refined MLIP | 359.67 | 125.23 | 54.76 | 203.06 |
@@
FigureΒ 7 shows the temperature-dependent properties calculated using the unrefined and refined MLIPs for UN, specifically (a) lattice parameter, (b) heat capacity, and (c) bulk modulus. Experimental values are added for reference. The lattice parameters computed using the unrefined MLIP are underestimated relative to the experimental values in the temperature range plotted from 300 to 2500 K (FigureΒ 7(a)). Kocevski et al. also report underestimated lattice parameters for UN from DFT-MD simulations,[31] demonstrating that the disagreement between the unrefined MLIP and experiment arises primarily from the DFT training dataset. Refinement improves agreement of lattice parameter with experimental data over the entire temperature range, reducing the degree of underestimation.
In FigureΒ 7(b), the experimental heat capacity data deflects to larger values at around 1600 K. Similar behavior is predicted by both the unrefined and refined MLIPs. The refined MLIP shows an improvement over the unrefined MLIP at temperatures above 2000 K with respect to experimental data.
@@
FigureΒ 7(c) shows that the refined MLIP for UN predicts lower values for the bulk modulus over the entire temperature range. The decrease in bulk modulus is consistent with weaker interatomic forces. Considering the significant improvement in bulk modulus at 0 K shown in TableΒ 5, a lower bulk modulus at higher temperatures is also likely a desirable result. Unfortunately, experimental bulk modulus data are not available at temperatures above 300 K for validation.
TableΒ 6 reports the defect energies for the unrefined MLIP, the refined MLIP, and DFT. The defect energies for both the unrefined and refined MLIP are reasonable compared to DFT, with the refined MLIP predicting slightly better Frenkel pair energies. Therefore, although the refined MLIP is re-trained to shifted atomic energies, the refinement process does not appear to negatively impact the defect energies.
@@
Certain physical properties are extremely sensitive to the hyperparameters used for the pre-trained MLIP. If the hyperparameters are not chosen judiciously, the refined MLIP may inherit some deficiencies from the unrefined MLIP that are only exacerbated during the refinement process. Since, the refinement process does not retrain to atomic forces, an important hyperparameter is the energy-to-force weight ratio in the loss function for pre-training. For this purpose, we developed another UN MLIP pre-trained with an energy-to-force weight ratio of 1:100 in the loss function, instead of the 1:1 ratio used in previous work,[29] but keeping all other hyperparameters the same. Subsequently, we repeat the refinement process and compare the 1:100 refined MLIP with the 1:1 refined MLIP. Compared to the 1:1 refined MLIP, the 1:100 refined MLIP has a more accurate thermal expansion coefficient across the temperature range and a more stable bulk modulus with respect to refinement generations (see Figure S15). However, certain defect energies and elastic constants are worse for the 1:100 refined MLIP compared to the 1:1 refined MLIP (see Tables S5 and S6). Although these preliminary results suggest that pre-training with a larger force weight can be beneficial for subsequent EXAFS refinement, more research is needed to determine whether this observation applies to other systems.

### v2

@@ Results
To achieve beyond DFT-level accuracy, we refine the pre-trained MLIP of uranium dioxide and uranium mononitride to match target synthetic EXAFS spectra. The unrefined MLIPs are pre-trained to DFT/DFT+U total energy and atomic forces, while the refined MLIPs are re-trained with transfer learning to the shifted atomic energies. Although only atomic-scale quantities are involved in the initial training and refinement process, the refined MLIP predicts improved macroscopic properties beyond EXAFS, that are not explicitly targeted in the training process. To validate the improvement and robustness of the refined UO2 and UN MLIPs, we calculate some relevant properties and compare the MLIP predictions with available experimental data, for temperature-dependent thermophysical properties (lattice parameters, bulk modulus and heat capacity), crystal properties (elastic constants and phonon dispersion), energetic properties (point defect energies), and transport properties (diffusion coefficients).
We first verify the accuracy of the UO2 and UN unrefined MLIPs pre-trained to the respective DFT+U and DFT datasets. Table 1 reports the energy-per-atom and force root-mean-square error (RMSE) between the unrefined MLIP-predicted values and the corresponding DFT+U/DFT reference values for UO2 and UN. The energy RMSE values are similar for both MLIPs and less than 10 meV/atom. The force RMSE for the UO2 MLIP is approximately 0.435 eV/Å, while the force RMSE for the UN MLIP is around 0.136 eV/Å. The relatively larger force RMSE value of UO2 is a consequence of the reduced DFT+U dataset consisting of more challenging atomic configurations, with fewer near-equilibrium low-energy configurations. For comparison, the RMSEs for UN (3.9 meV/atom for energy and 0.136 eV/Å for forces) are only slightly higher than those reported previously (2.48 meV/atom for energy and 0.113 eV/Å for forces) for the same HIP-NN architecture trained on the same dataset [1]. Furthermore, the RMSEs for both UO2 and UN are of the same magnitude to the reported values for a different MLIP architecture (4.5 meV/atom for energy and 0.380 eV/Å for forces for UO2 and 2.3 meV/atom for energy and 0.072 eV/Å for forces for UN), demonstrating that our model achieves the expected level of accuracy on these training datasets [58, 1].
| Material | Energy (meV/atom) | Force (eV/Å) |
| UO2 | 6.5 | 0.435 |
| UN | 3.9 | 0.136 |
Because the refined MLIPs are trained to shifted atomic energies, the RMSEs for the refined MLIPs with respect to the DFT+U/DFT energies are significantly higher (compare Figures S9(a) and S10(a) with Figures S9(c) and S10(c)). However, the energy errors are almost entirely positive for UO2 and negative for UN, demonstrating that the energy shifts are systematic and internally consistent. Furthermore, although atomic forces were not utilized during retraining, the refined MLIPs predict the DFT+U/DFT forces with nearly the same accuracy as the unrefined MLIPs (compare Figures S9(b) and S10(b) with Figures S9(d) and S10(d)). Some increase in force errors is expected, as the DFT+U/DFT-level forces are not consistent with the shifted atomic energies. However, the relatively small increase in force errors demonstrates that the refined MLIPs did not completely “forget” the fundamental physics contained in the DFT+U/DFT training datasets.
@@
Since the refinement process is performed with respect to the synthetic target EXAFS spectra for both U and O, we verify that the refined MLIP predicts spectra in closer agreement with the experimental EXAFS spectrum for U. Because refinement is performed at a single temperature of 75 K, we also verify the improvement for EXAFS spectrum prediction at 300 K.
@@
Figure 2 presents a comparison of the experimental EXAFS spectrum for uranium in UO2 with those computed using unrefined and refined MLIPs at 77 K (a) and at 300 K (b). At both temperatures, the spectral peaks shift towards higher k-values, aligning more closely with the experimental data. At 77 K, the peak amplitude of the refined MLIP increases slightly compared to the unrefined MLIP. However, both the unrefined and refined spectra have an enhanced peak amplitude compared to the experimental spectra. At 300 K, we observe better agreement of the peak amplitude with the experimental spectrum.
@@
Even though the amplitudes for the synthetic perfect crystal spectrum are significantly larger than the corresponding ones in the experimental spectrum (see Figure S6), the refined MLIP spectral peak amplitudes are in reasonable agreement with experiment. Overfitting to the synthetic spectra could induce unusually strong interatomic forces, reducing thermal vibrations and producing overly large peak amplitudes. Careful transfer learning (see Simulation Details) helps avoid overfitting and catastrophic forgetting of the initial DFT+U training dataset. Figure S11 also shows systematic and gradual improvement in the MLIP-generated EXAFS spectra with respect to the synthetic target spectra, suggesting that the EXAFS refinement approach converges smoothly.
@@
Table 2 reports the UO2 elastic constants at 296 K for both MLIPs (unrefined and refined) compared to the experimental values at 296 K [16]. The refined MLIP shows a significant improvement in the underestimated and constants, while still overestimating the constant.
| Elastic constants | (GPa) | (GPa) | (GPa) |
| Experiment | 389.3 1.7 | 118.7 1.7 | 59.7 0.3 |
| Unrefined MLIP | 346.02 | 140.18 | 44.68 |
| Refined MLIP | 367.10 | 143.82 | 49.20 |
Figure 3 shows the predicted temperature-dependent (a) lattice parameter, (b) heat capacity, and (c) bulk modulus using the unrefined and refined MLIPs in comparison to experimental data. The systematic and gradual improvement of these properties at each refinement iteration is shown in Figure S12. As shown in Figure 3(a) and (c), the most significant improvements achieved through MLIP refinement are the enhanced accuracy of lattice parameter and bulk modulus with respect to temperature. The percent improvements in lattice parameter and bulk modulus of approximately 1.5% and 15%, respectively, are comparable with those achieved when refining a DFT-trained MLIP for titanium using differentiable trajectory reweighting [Röcken2024].
@@
Because EXAFS spectra depend strongly on the local atomic environment and the interatomic distances around the absorbing atom, the improvement in lattice parameter can be attributed to a better agreement with the target EXAFS spectra, which was generated with the experimental lattice parameter. However, the MLIP refinement process does not train directly to lattice parameter data. Thus, the lattice parameter temperature dependence represents an out-of-sample test case. Remarkably, the refined MLIP produces improved lattice parameters over a wide range of temperatures from 300 K to 3000 K, even though the refinement is performed at a single low temperature (75 K). Previous work also reports similar improvement in lattice parameter over a wide range of temperatures when refining with experimental data at a single temperature using differentiable trajectory reweighting [Röcken2024]. Figure 3(a) demonstrates that the steady overprediction of lattice parameter given by the unrefined MLIP gradually deflects above 2000 K. This deflection is also improved in the lattice parameters predicted by the refined MLIP, suggesting that the refined MLIP is more stable at higher temperatures. This improvement can be more clearly observed from the thermal expansion plot at higher temperatures, from 2500 K to 3000 K shown in Figure S13. Such consistent improvement over a range of temperature is not observed when the re-training is done without freezing certain HIP-NN layers (Figure S14 and Figure S15), which implies that freezing some layers during transfer learning eliminates possible overfitting to the small refinement dataset.
@@
Figure 3(b) shows the temperature-dependent heat capacity for unrefined and refined MLIPs compared to the experimental results. The increase in heat capacity with temperature for UO2 has been attributed to different factors in different temperature ranges. At low temperatures up to 1000 K, the strongest contribution to heat capacity are harmonic lattice vibrations, however between 1000 K to 1500 K, there is an increase in anharmonicity. The sharp increase of heat capacity observed from 1500 K up to 2670 K, has been linked to the formation of lattice and electronic defects [28]. An abrupt transition in heat capacity, named Bredig temperature is observed experimentally at 2670 K and is considered a result of the saturation of defect concentration [internationalsurl'étudedestransformationscristallinesàhautetempératureau-dessusde2000K1972]. While both refined and unrefined MLIPs reproduce this transition phenomenon, the refined MLIP predicts a slightly more accurate transition temperature compared to the unrefined MLIP, again indicative of a better description of UO2 at high temperatures.
@@
Figure 3(c) shows improvement in the refined MLIP predictions for bulk modulus as a function of temperature. The predicted bulk modulus using the unrefined MLIP is underestimated with respect to the experimental data of Hutchings et al. [27], except at lower temperatures, where the MLIP matches the experimental value at 296 K. In contrast, the refined MLIP predictions are consistent over the entire temperature range and in closer agreement–even at high temperatures–with the experimental values. Overall, the refined MLIP predicts higher bulk modulus compared to the unrefined MLIP, consistent with higher EXAFS peak amplitudes and a stiffer response to volumetric compression.
@@
Figure 4 shows that the phonon dispersion spectra are similar for both the unrefined and refined MLIP. This similarity reflects that UO2 lattice vibrations are not affected significantly after refinement. Since, the synthetic target spectra is computed for a perfect crystal, overfitting to the synthetic target spectra during refinement would lead to stiffer vibrations and degrade phonon dispersions. However, the phonon spectra remain relatively unchanged, demonstrating that the refinement process improves the structural properties while not causing the refined MLIP to forget lattice vibration contributions contained in the DFT+U original training dataset.
@@
Nuclear fuels are continuously exposed to neutron irradiation, leading to the formation of vacancies, interstitial and dislocation defects. Defect energies are critical properties, as they can influence mechanical degradation and ultimately fuel operation lifetime. The unrefined MLIP yields point defect energies that align well with DFT+U results [11, 64]. Since the refinement process shifts the atomic energies in the MLIP, it is essential to ensure that point defect predictions remain accurate. Table 3 reports the point defect energies obtained from both the unrefined and refined MLIPs, compared to experimental and DFT+U literature values [11, 64, 40]. The Schottky defect energy predicted by both MLIPs falls within the experimental range. Similarly the uranium and oxygen Frenkel pair defect energies are within the DFT+U range, although they are slightly overestimated relative to experimental data. Considering the broad uncertainty in both experimental values and DFT+U calculations, the point defect energies calculated from both the unrefined and refined MLIPs are considered reasonable. This confirms that the refinement process does not negatively affect defect energies.
Subsequent validation of transport properties is important when training to only static structural data, such as EXAFS spectra. For this reason, we compare the diffusion coefficient for oxygen in UO2 predicted with the unrefined and refined MLIPs. The simulation methods used to calculate these values are provided in the Simulation Details section and Supporting Information (see Figure S16). Figure 5 provides the oxygen diffusion coefficients as a function of inverse temperature from 2100 K to 3000 K. The refinement process increases the diffusion coefficient values approximately by an order of magnitude. For better comparison with experimental data, the activation energies for oxygen diffusion in UO2 are calculated using an Arrhenius fit to the diffusion coefficient. Figure 5 shows that two distinct linear regimes exist. However, the transition temperature that separates these low and high regimes differ between the unrefined MLIP and refined MLIP by 100 K, with the refined temperature of 2700 K matching better the experimental Bredig temperature of 2670 K.
Table 4 reports the activation energy for low- and high-temperature regimes obtained from oxygen diffusion in UO2 and calculated with the unrefined and refined MLIPs. In the high-temperature regime, the activation energy agrees closely with the experimental values [34]. The activation energy in the low-temperature regime is overestimated for both unrefined and refined MLIPs. Notably, the re-training has minimal impact on the activation energy in both ranges of temperature, demonstrating that the MLIP has retained valuable transport information from the DFT+U dataset.
@@
To further validate our methodology, we refine an MLIP for uranium mononitride that was pre-trained to DFT data [1]. In contrast to UO2, DFT+U calculations were not required to properly describe the magnetic and metallic character of UN. Therefore, the unrefined UN MLIP predictions do not differ as strongly from experiment, especially for the lattice parameter. Consequently, unlike UO2, the EXAFS peak positions obtained from the unrefined MLIP already align well with the synthetic target EXAFS spectra for UN (see Figure S17). Although less improvement is needed for UN system than for UO2, this test case demonstrates how it is still important to carefully refine MLIPs without depreciating their accuracy for other properties that are already predicted well from the DFT training dataset.
@@
As in the previous UO2 case, we use synthetic target spectra for U and N at a single low temperature of 50 K for the refinement procedure. Once again, we verify that the agreement with synthetic spectra also improves agreement with the experimental U EXAFS spectra, even at higher temperatures. In Figure 6, we compare the U EXAFS spectra for unrefined and refined MLIPs with respect to experimental EXAFS spectra at (a) 80 K and (b) 300 K. At a high level, the unrefined and refined MLIPs reproduce almost every broad oscillation present in the experimental EXAFS spectra. The EXAFS peak positions of the refined MLIP align slightly better with experimental data compared to the unrefined MLIP, as evidenced by the moderately lower RMSE value. However, the simulated spectra for both the unrefined and refined MLIPs are smooth and do not capture certain sharp features/peaks found in the experimental spectra. We attribute this smoothness to the cutoff radius of 6 Å used to calculate the simulated spectra. Increasing the cutoff can capture more nuanced structural information, resulting in more detailed EXAFS spectra [47, 62]. However, a greater cutoff radius increases the number of scattering paths from distant coordination shells, which dramatically increases the computational cost of calculating the EXAFS spectra. Selecting a 6 Å cutoff radius is adequate for resolving key structural features in EXAFS while keeping computational costs manageable during the refinement process.
@@
As in the UO2 case, the synthetic target EXAFS spectrum of uranium in UN has a larger amplitude compared to the experimental spectra. The synthetic spectra are calculated for a perfect crystal structure, which eliminates the dampening of peaks due to atomic vibrations. Despite the larger amplitude of target spectra, the refined MLIP actually predicts a slight reduction in peak amplitude (see Figure S17), which improves agreement with the experimental EXAFS spectrum (Figure 6). The refinement also causes a slight leftward shift in both U and N EXAFS spectra, mainly in the range of the wavenumber k = 8 to 15. The leftward shift of the EXAFS peaks leads to an increase in the interatomic distances, signifying a decrease in the interatomic forces. With more flexible vibrations, the mean-square-displacement of UN atoms increases, which slightly dampens the spectra. The evidence of weaker interatomic forces is supported by certain properties used to validate the refined MLIP, discussed below.
@@
To validate the refined MLIP for UN against structural properties, Table 5 reports the elastic constants at 290 K using the unrefined and refined MLIPs as well as the experimental values at 290 K [50]. The values for and show an improvement, while the value is underestimated after refinement.
| Elastic constants | (GPa) | (GPa) | (GPa) |
| Experiment | 423.9 0.6 | 98.1 0.9 | 75.7 0.2 |
| Unrefined MLIP | 381.59 | 123.30 | 47.65 |
| Refined MLIP | 351.47 | 116.35 | 51.33 |
@@
The temperature-dependent properties calculated using the unrefined and refined MLIPs for UN in the temperature range between 300 K to 2500 K are shown in Figure 7. Specifically, we calculate: (a) lattice parameter, (b) heat capacity, and (c) bulk modulus. Available experimental values are added for reference. The lattice parameters (Figure 7(a)) calculated using the unrefined MLIP are underestimated relative to the experimental values [22] across the temperature range. Kocevski et al. also report underestimated lattice parameters of 0.5% for UN from DFT-MD simulations [31], demonstrating that the disagreement between the unrefined MLIP and experiment arises primarily from the DFT training dataset. Refinement reduces the underestimation of lattice parameter with respect to experimental data over the entire temperature range, with a similar degree of improvement to previous work using differentiable trajectory reweighting [Röcken2024].
@@
Similar trends for heat capacity are predicted by both the unrefined and refined MLIPs as shown in Figure 7(b). The refined MLIP shows some improvement over the unrefined MLIP at temperatures above 2000 K with respect to experimental data. The bulk modulus as function of temperature calculated with the refined MLIP in Figure 7(c) is lower in comparison to the unrefined MLIP over the entire temperature range. The decrease in bulk modulus is consistent with weaker interatomic forces. However, the refinement process appears to have slightly over-corrected the bulk modulus at low temperatures as the refined MLIP underpredicts the experimental data point. Unfortunately, experimental bulk modulus data are not available at temperatures above 300 K for further validation.
Finally, we calculate point defect energies for UN and report them in Table 6. We compare the unrefined and refined MLIP values with DFT. Both MLIPs predict point defect energies that are in agreement with DFT, yet the refined MLIP shows a slight improvement for the uranium and nitrogen Frenkel pair defects compared to the predictions by the unrefined MLIP. This confirms again that despite the being re-trained to shifted atomic energies, the refinement process does not adversely affect the defect energies.
@@
Certain physical properties, in particular the bulk modulus, are extremely sensitive to the hyperparameters used for the pre-trained MLIP. If the hyperparameters are not chosen judiciously, the refined MLIP may inherit some deficiencies from the unrefined MLIP that are only exacerbated during the refinement process. Since, the refinement process does not retrain to atomic forces, an important hyperparameter is the energy-to-force weight ratio in the loss function for pre-training. For this purpose, we developed another UN MLIP pre-trained with an energy-to-force weight ratio of 1:100 in the loss function, instead of the 1:1 ratio used in previous work [1], but keeping all other hyperparameters the same. Subsequently, we repeat the refinement process and compare the 1:100 refined MLIP with the 1:1 refined MLIP. Compared to the 1:1 refined MLIP, the 1:100 refined MLIP has a more accurate thermal expansion coefficient across the temperature range and a more stable bulk modulus with respect to refinement generations (see Figure S18). However, certain defect energies and elastic constants are worse for the 1:100 refined MLIP compared to the 1:1 refined MLIP (see Tables S5 and S6). Although these preliminary results suggest that pre-training with a larger force weight can be beneficial for subsequent EXAFS refinement, more research is needed to determine whether this observation applies to other systems.